Given x=-2 and y=3, find the value of the expression x^2y^3. Thanks for your help.
Found 2 solutions by fion6002, jim_thompson5910:
Answer by fion6002(19)

x^2y^3
= (-2)^2*(3)^3
= 4*27
= 108

Answer by jim_thompson5910(35256)

Replace x with -2 and y with 3 to get...

x^2y^3 = (-2)^2*(3)^3 = 4*27 = 108

So, x^2y^3 = 108 when x = -2 and y = 3.